5t^-10t=0 solve by factoring
Answer by checkley77(12844)

5t^[2]-10t=0 I'll assume you forgot the (^2)
t(5t-10)=0
5t-10=0
5t=10
t=10/5
t=2 ans.
Proof:
5*2^2-10*2=0
5*4-20=0
20-20=0
0=0